In 2019, ESERAH purchased property with mineral resources for ₱28,000,000. The property had a residual value of ₱5,000,000. However, ESERAH is required to restore the property to the original condition at a discounted amount of ₱2,000,000.

In 2019, ESERAH spent ₱1,000,000 in development cost and ₱3,000,000 in building. In 2021, an amount of ₱4,000,000. Was spent for additional development on the mine.

Production began in 2020 and the tons extracted totaled 3,000,000 in 2020 and 2,500,000 in 2021. The estimated remaining tons to be extracted totaled 7,000,000 and 3,500,000 at the end of 2020 and 2021, respectively.

Questions:

Based on the above data, answer the following:

1. What is the depletion base?

2. What amount of depletion should be recognized in 2020?

3. What amount of depletion should be recognized in 2021?

4. What is the carrying amount of the mineral property on December 31,2021?



1
Expert's answer
2021-06-16T13:31:46-0400



a) Depletion base = Cost of natural resources= 28,000,000

b)Depletion in 2020=3,000,000 *3.00=9,000,000

c) Depletion in 2021 = 2,500,000* 3.50=8,750,000

d) Carrying amount =21,000,000